### v1

@@ III.1 Quantum Refrigerator
the refrigeration mode (R) is predominant when the thermal gap between the two reservoirs is small Fig.3. We will examine two values of $\alpha$: the first promotes long-range correlations ($\alpha=0.25$), and the second favors short-range correlations ($\alpha=1.2$) . We will also investigate the behavior of thermodynamic quantities such as efficacy, $\frac{Q_{c}}{N}$, and the scaling factor per spin, $\frac{\Pi}{N}$
We notice in Fig.7($a$-$b$) that $Q_{c}/N$ for both values of $\alpha$ is of the same magnitude , for $\alpha=0.25$, small chains perform slightly better at low fields . We observe that it peaks just before the critical point $h_{i}=1$ and then slightly declines before and after the phase transition. It stabilizes after $h_{i}=2$ The contour lines of $\frac{Q_{c}}{N}$ for small $h_{i}$ are very tightly packed, particularly for short chains, showing greater variability and a more significant decrease compared to longer chains, which have less dense contour lines. It is observed that the contour lines indicate a slower and more stable decrease for all $N$ after $h_{i}=1.5$. In the case of $\alpha=1.2$, long chains outperform short ones before and around the critical field. In the critical field, $h_{\text{critique}}=0.7$, we observe a pronounced drop, additionally, we note that short-range correlations lead to a steeper and more abrupt decline compared to long-range correlations.We notice that for $Q_{c}/N$, the contour lines are less tight, with significant values for large $N$. Unlike the previous case, there is a more substantial decrease for long and intermediate chains, while the decrease is slower for short chains. A stabilization occurs for all chain sizes starting from $h_{i}=1.5$.
Figure 7: Contour plot of $Q_{c}$ and $\eta_{R}$ as a function of $h_{i}$ and N.
Figure 8: COP of the refrigerator versus $h_{i}$, for different system sizes $N$ across the critical point.
Figure 9: Scaling factor per spin $\Pi_{R}/N$ as a function of $h_{i}$ for different N.
The plot of COP as a function of $h_{i}$ Fig.8($a$) for different $N$ at $\alpha=0.25$ for $\frac{Q_{c}}{|Q_{c}+Q_{h}|}$ shows better efficiency for small chains (10 to 40), with a faster growth than larger chains, particularly near the critical field $h_{i}=1$. There is a distinct difference in behavior between large and small $N$, as in the post-critical region there is a sharp drop for small chains (10 to 40), with stabilization occurring from $h_{i}=1.8$. For intermediate and large chains, we observe a slight increase before the critical magnetic field, which becomes more pronounced after the phase transition, Starting from $h_{i}=1.5$, a shift in efficiency is observed: larger chains exhibit greater efficiency compared to intermediate and shorter chains. This highlights a stronger size dependence after the phase transition, particularly for higher values of $h_{i}$. Beyond the critical point, small chains struggle to recover their refrigeration efficiency, with $Q_{c}/N$ remaining low and showing no significant improvement as $h_{i}$ increases. This highlights the difficulty for smaller systems to adapt to the post-transition regime, resulting in a diminished capacity to extract heat effectively. In contrast, larger chains demonstrate a partial recovery in $Q_{c}/N$ as $h_{i}$ exceeds 1.5, suggesting their greater adaptability and resilience in the altered dynamics of the post-critical regime. While they do not reach the efficiency levels observed before the transition, large systems still maintain some level of refrigeration capability, indicating their robustness in maintaining performance despite the phase change.
The contour lines Fig.7($c$) show that they are low for small $h_{i}$, particularly for long chains. As $h_{i}$ increases (with intermediate magnetic fields $h_{i}$ from 0.5 to 1.5), the lines become concave and expand as $N$ decreases, centered around the critical field where they reach a maximum for short chains before dropping sharply after higher $h_{i}$. It is noted that for strong magnetic fields, the efficiency of long chains increases gradually.
The figure Fig.8($b$) showing the efficiency for $\alpha=1.2$ as a function of $h_{i}$ and $N$ demonstrates that short chains are more efficient than long chains, similarly to previous cases. For small $h_{i}$, there is a slight and similar increase for both long and short chains. As the critical field $h_{i}=0.7$ is approached, the slope increases, more pronounced for long chains, which rise rapidly to approach the efficiency of short chains for intermediate fields. For higher magnetic fields, there is a slower increase for long chains, with stabilization around $h_{i}=2$, and a faster increase for short chains. For large chains, saturation is observed for all values of $h_{i}$, and the increase in COP is more significant for short chains, indicating that they undergo a more gradual transition and can maintain effective refrigeration for all values of $h_{i}$.
The efficiency contour lines are less tight compared to the previous case, showing an increase in efficiency as $h_{i}$ rises for all $N$, with a leveling of the efficiency near the critical field $h_{i}=0.7$. Efficiency increases in the intermediate field range, and the contour lines become vertical. For stronger fields, starting from $h_{i}=1.7$, the lines curve and become more concave, less uniform, and more efficient for short chains than for long ones, with a peak for short chains ($N=10$ to $20$) at $h_{i}=2.5$ (see Fig.7($d$)).
The scaling factor per spin, $\Pi_{R}/N$, is influenced by the parameter $\alpha$, unlike $Q_{c}/N$, where the values are not of the same order. An increase in $\alpha$ amplifies the fluctuations in our quantum refrigerator. It is also important to note that increasing $\alpha$ brings the temperatures of the reservoirs closer together. For $\alpha=0.25$, there is a pronounced peak at low $h_{i}$ values, especially for very long chains, which diminishes as $h_{i}$ increases. This decline is more significant for shorter chains near the phase transition at $h_{i}=1$, where the slope decreases and stabilizes performance for larger $h_{i}$ values. At $h_{i}=2.25$, the performance of long and short chains converges, demonstrating a lack of size dependence for higher $h_{i}$ values.For $\alpha=1.2$, before the paramagnetic phase, performance stabilizes with a slight decline as $h_{i}$ increases across all chains. At the critical field $h_{i}=0.7$, there is a sharp drop in the scaling factor, signaling underperformance and instability, particularly for longer chains, which remain more efficient than smaller $N$ but exhibit a steeper decline. As $h_{i}$ increases, there is a modest decrease followed by stabilization (see Fig.(9)).
Overall, the system’s behavior strongly depends on the chain size and the correlation regime, with each configuration offering specific advantages depending on the magnetic field. The performance per spin is better in the ferromagnetic phase, with efficiency varying depending on the type of correlation $\alpha$.
@@ III.2 Quantum Heat Engine
We will see in this section the evolution of the work per spin ($W/N$) and efficiency ($\eta$) as a function of the initial transverse field ($h_{i}$) for different values of $\alpha$, $\beta_{h}$, and $\beta_{c}$, analyzing the effects of the phase transition and the system’s behavior for small and large chains ($N$). Each graph in Fig.(10) highlights how interactions and thermal conditions influence the system’s performance.
Figure 10: Work per spin $W/N$ as a function of $h_{i}$ for different N.
In Fig.10($a$), with $\alpha=0.9$, $\beta_{h}=0.5$, and $\beta_{c}=1.0$, it is observed that for large values of $N$ (e.g., $N>70$), the work per spin is higher in the region where $h_{i}\leq 1.0$. Beyond this point, the slope decreases, leading to a stabilization of the curve. The maximum work per spin, $W/N$, occurs around $h_{i}=0.25$, indicating a more effective capture of short-range correlations near the critical point. It is noteworthy that, across all regions, long chains exhibit a higher $W/N$ compared to short chains. After the critical point, $W/N$ declines rapidly, especially for large systems, signaling a significant shift associated with the phase transition. In contrast, for smaller systems, the maximum is less pronounced, and the work is lower, highlighting the dominance of thermal fluctuations and local correlations. When comparing the sizes of the chains, $W/N$ remains higher for long chains throughout the $h_{i}$ range from 0 to 2.5, demonstrating greater efficiency in capturing short-range correlations for longer chains. In Fig.10($b$), corresponding to $\alpha=1.2$, $\beta_{h}=0.2$, and $\beta_{c}=0.9$, the work per spin is similarly elevated for larger values of $N$, peaking around $h_{i}=0.35$, which aligns with the estimated critical point for $\alpha>1$ at $h=0.7$. Long chains are more adept at capturing short-range correlations than their shorter counterparts prior to the transition, thereby accounting for a higher $W/N$. After $h_{i}=0.7$, $W/N$ experiences a steep decline, occurring more rapidly than in the previous case, while the work per spin of long chains remains superior to that of short chains. This observation indicates that longer chains continue to effectively capture short-range correlations even subsequent to the phase transition. In contrast, shorter chains exhibit a more stable behavior, suggesting reduced sensitivity to fluctuations in $h_{i}$. The Fig.10($c$), with $\alpha=0.55$, $\beta_{h}=0.2$, and $\beta_{c}=0.667$, shorter chains ($N<30$) exhibit a higher $W/N$ for $h_{i}$ ranging from 0 up to the critical point at $h=1.0$, where a transition occurs, allowing long chains to become more efficient following the phase transition; this alteration does not apply to short-range correlations. The peak of $W/N$ is observed around $h_{i}=0.25$, indicating that smaller chains are more effective in capturing intermediate-range correlations prior to the phase transition. However, after $h_{i}=1$, longer chains ($N>70$) surpass shorter chains in terms of $W/N$, reflecting a superior ability to capture long-range correlations. This implies that for $\alpha=0.55$, the phase transition significantly alters the behavior of long chains post $h_{i}=1$, while shorter chains maintain their efficiency before this transition. Fig.10($d$), with $\alpha=0.2$, $\beta_{h}=0.6$, and $\beta_{c}=2.0$, exhibits behavior akin to the previous case, with the performance of long chains surpassing that of shorter chains before the critical point at $h=1$, particularly near $h=0.4$. For shorter chains ($N<30$), $W/N$ is elevated for $h_{i}\leq 0.4$, peaking around $h_{i}=0.8$ for longer chains. Beyond this threshold, $W/N$ declines rapidly, indicating a loss of efficiency in the smaller chains. In contrast, after the critical field, starting from $h_{i}=1$, there is a significant drop in $W/N$. This illustrates that long-range correlations in long chains at the phase transition accelerate the decline of $W/N$.
The contour lines of the work per spin $W/N$ presented in Fig.11 show behavior that depends on $\alpha$ and the thermal gap. When $\alpha$ is small and favors long-range correlations, there are two peaks: one for short chains and the other for long chains unlike $\alpha=0.25$, we observe a peak before the transition (for small $N$) and another at the critical field (for large $N$). When $\alpha=0.55$, there is a peak before the critical field and another after the phase transition, which is farther apart with a lower $W/N$ at the critical field compared to the previous case. When short-range correlations are favored ($\alpha=0.9$ and $1.2$), we observe a single peak before the critical field for long chains. The only difference between the two cases lies in the variability and a smaller decrease in $W/N$ before the phase transition for $\alpha=1.2$, with more vertical contour lines. This detailed analysis presents the behavior of the efficiency $W/Q_{h}$ and the work per spin $W/N$ as a function of $h_{i}$ for different values of $\alpha$, $\beta_{h}$, and $\beta_{c}$.
Figure 11: Contour plot of Work per spin W/N and efficient $\eta$ as a function of $h_{i}$ and N.
With $\alpha=0.9$, $\beta_{h}=0.5$, and $\beta_{c}=1.0$, the efficiency $W/Q_{h}$ shows a gradual decline as $h_{i}$ increases. The phase transition, occurring at $h=1$, corresponds to the critical point $h_{c}$ for $\alpha\leq 1$. Prior to this transition, longer chains (with lower $N$) display greater efficiency compared to shorter chains. However, near the critical point, the efficiency of longer chains aligns closely with that of shorter ones. Beyond $h=1.25$, the disparity between short and long chains becomes more apparent, accompanied by a marked decrease in efficiency for both. In the second figure, with $\alpha=1.2$, $\beta_{h}=0.2$, and $\beta_{c}=0.9$, we observe a pattern reminiscent of the case with $\alpha=0.9$, where short-range correlations are predominant. The critical field is located at $h_{i}=0.7$. Prior to the phase transition, there is a slight reduction in efficiency for chains of all sizes, with large $N$ chains exhibiting greater efficiency. At the critical point, a distinctive behavior emerges with a sharp drop in efficiency, resulting in a significant decline across all chains, leading to an identical efficiency independent of chain length in the post-critical region. Efficiency remains relatively high for small values of $h_{i}$ (between 0 and 0.5), reaching a peak at very low fields. Large chains (with higher $N$) demonstrate superior efficiency compared to smaller chains across the entire range of $h_{i}$, both before and after the critical field. This stability in larger chains following the transition highlights a regime where short-range correlations prevail, maintaining efficiency despite variations in the transverse field $h_{i}$. In the third figure, with $\alpha=0.55$, $\beta_{h}=0.2$, and $\beta_{c}=0.667$, which favors intermediate range correlations, the efficiency $W/Q_{h}$ exhibits a gradual decline as $h_{i}$ increases, with nearly identical efficiency for large $N$ values, indicating pre-saturation for long chains in the range $h_{i}=0$ to 0.5, just before the phase transition. Around the critical field $h_{c}=1$ for $\alpha\leq 1$, a stabilization is noticeable, where efficiency remains almost unchanged. Beyond $h_{i}=1.5$, a slight reduction in efficiency is observed for all chain lengths. The curve’s behavior suggests heightened sensitivity of the system to changes in $h_{i}$ near the critical point. The Fig.12($a$), with $\alpha=0.2$, $\beta_{h}=0.6$, and $\beta_{c}=2.0$, the efficiency is initially high but steadily declines as $h_{i}$ increases, reaching the critical point at $h_{i}=1$. This behavior reflects a phase transition, where the system’s sensitivity to the transverse field $h_{i}$ is notably heightened for small values of $\alpha$. The slope increases and the decline becomes more rapid for larger $h_{i}$, emphasizing the influence of long-range correlations in stronger fields. Post-transition, shorter chains demonstrate greater stability with fewer fluctuations compared to longer chains. For large $N$, the efficiency becomes more uniform, exhibiting faster saturation, more pronounced than in the case of $\alpha=0.55$, though this saturation occurs in the post-transition region, accentuated by long-range correlation effects. In stronger fields, such as $h_{i}=2$, intermediate-length chains show superior efficiency, indicating a shift in the system’s operational regime where long-range correlations become less effective for longer chains (see Fig.12).
Figure 12: Efficiency $\eta$ as a function of $h_{i}$ for different $N$.
The contour lines of efficiency $\eta$ show a dependence on $\alpha$. Similar to $W/N$, there are two distinct behaviors. For small $\alpha$, we observe a peak for long chains at very low $h_{i}$, with the contour lines becoming more horizontal for medium and large $N$, then more vertical for long chains, which is less noticeable for $\alpha=0.55$. For larger $\alpha$, there is a clear peak for medium and large chains at low $h_{i}$, with vertical contour lines where efficiency increases horizontally. This effect is more pronounced for $\alpha=1.2$, which shows slower variability compared to $\alpha=0.9$ (Fig.11(e)-(h)).
Unlike work, the efficiency for all the parameters previously mentioned is dependent on $N$. To increase efficiency for fixed parameters, it is sufficient to increase the number of particles, which is not always the case for $W/N$. We observe that the efficiency ($\eta$) and the work per spin ($W/N$) exhibit different behaviors, as the work per spin is not proportional to the chain size for small and intermediate values of $\alpha$. In contrast, the efficiency depends directly on the size of the chain: the larger $N$ becomes, the greater the efficiency ($\eta$). This discrepancy can be explained by the nature of the correlations and fluctuations induced by the temperature difference between the reservoirs. For larger values of $\alpha$ (e.g., 0.9 and 1.2), we observe that when the temperature difference ($\Delta T$) is significant (as with $\alpha=1.2$), the system achieves a higher work per spin ($W/N$) compared to $\alpha=0.9$, but with a lower efficiency. This can be attributed to thermal fluctuations and agitation, which are reduced when $\Delta T$ is smaller. The same interpretation applies to long-range and intermediate-range correlations, where the effect of temperature fluctuations diminishes with a smaller temperature gradient, leading to different performances in terms of both work and efficiency. Thus, the relationship between efficiency and work per spin reflects the complex interplay of thermal fluctuations and correlations in the system, especially across different interaction ranges.
III.3 Scaling factor of the quatum heat engine
The scaling factor $\Pi=\frac{W}{\eta_{\text{Carnot}}-\eta}$ measures how effectively an engine produces work despite having an efficiency lower than the Carnot limit. A high scaling factor means that the system is optimized to extract significant work, even in non-ideal conditions. In quantum engines, a higher $\Pi$ reflects more efficient energy conversion through quantum effects like correlations and superposition. It also suggests lower energy losses, allowing the system to approach Carnot efficiency. Overall, a high scaling factor indicates that the system is productive, efficient, and resilient to fluctuations.
To observe the impact and influence of correlations without considering the size of the chain, we will normalize by dividing by $N$.We will discussed this in more detail in the appendix Appendix D. The corrected interpretation of the provided figures focuses on the scaling factor $\Pi$, defined as
@@ III.3 Scaling factor of the quatum heat engine
where $W$ is the work done by the system, $\eta_{\text{Carnot}}=1-\frac{\beta_{h}}{\beta_{c}}$ is the Carnot efficiency, and $N$ is the number of particles in the system . The figures 13 illustrate $\Pi/N$ as a function of the transverse magnetic field $h_{i}$, highlighting the behavior of the system for varying values of $N$, ranging from 10 to 100, for different configurations of $\alpha$, $\beta_{h}$, and $\beta_{c}$. Fig.13($a$) $\alpha=0.9$, $\beta_{h}=0.5$ and $\beta_{c}=1.0$, $\Pi/N$ starts at high values for small $h_{i}$ and decreases significantly near the critical magnetic field $h_{i}=1$, indicating a strong phase transition. Large chains ($N>70$) show a sharp decrease in $\Pi/N$, while small chains show a more gradual decline. After $h_{i}=1$, $\Pi/N$ for large chains remains stable but lower compared to small $h_{i}$ values, indicating a more consistent capture of long-range effects. $\Pi/N$ drops significantly after the transition, reflecting their inability to maintain high performance in the long-range regime. In , corresponding to $\alpha=1.2$, Fig.13($b$)$\beta_{h}=0.2$ and $\beta_{c}=0.9$, $\Pi/N$ is high for large values of $N$ (for example, $N>70$) in the region where $h_{i}\leq 0.7$. The maximum occurs around $h_{i}=0.3$, reflecting better adaptation to critical fluctuations near the estimated transition point for $\alpha>1$. Beyond the critical point $h_{i}=0.7$, there is a sharp drop in the system’s performance, with a less steep slope for small chains. After $h_{i}=1$, $\Pi/N$ decreases for large chains but remains higher compared to small chains ($N<30$), indicating that large chains are more effective in maintaining high performance before the transition but experience a more intense drop in performance than small chains, where long-range correlations dominate. Large chains show a rapid decrease in $\Pi/N$, reflecting their limited capacity to harness critical effects. At $\alpha=0.55$ Fig.13($c$), $\beta_{h}=0.2$ and $\beta_{c}=0.667$, shows that $\Pi/N$ is initially high, especially for small chains. Where long-range correlations are moderate before the critical field $h=1$, in the ferromagnetic phase, they force the particles to align better than in long chains, thus ensuring better $\Pi/N$ performance. After the transition phase $h_{i}=1$, large chains dominate in terms of $\Pi/N$, reflecting their superior performance in exploiting intermediate correlations in the ferromagnetic phase at higher field strength. This behavior indicates that large chains are more adaptable and efficient after the transition, whereas small chains are more efficient in the ferromagnetic phase, before the phase transition. In Fig.13($d$), with $\alpha=0.2$, $\beta_{h}=0.6$ and $\beta_{c}=2.0$, where the interactions are the longest, we observe that before $h_{i}=0.25$, the short chains perform better, and there is a shift at this point towards the dominance of long chains. $\Pi/N$ peaks at $h_{i}=0.6$ before the critical point but decreases rapidly beyond this point, showing poor performance in the post-transition regime. We also observe that $\Pi/N$ decreases slowly after $h_{i}=1$, indicating an effective capture of long-range correlations. This highlights the dominance of long-range effects for large chains before the phase transition. It is also important to note that, as in the previous section, there is a slight shift between long and short chains post-phase transition, as the short chains stabilize more quickly because the fluctuations and thermal perturbations, amplified by long-range correlations in the paramagnetic phase.
Figure 13: Contour plot of the scaling factor per spin $\Pi_{R}/N$ as a function of $h_{i}$ and N.
The impact of $\alpha$, $\beta_{h}$, and $\beta_{c}$ on the system is significant. For high $\alpha$ values (e.g., $\alpha=1.2$), short-range correlations dominate, leading to a smoother transition and more stable $\Pi/N$. Large chains benefit significantly from short-range interactions before $h_{c}$. For low $\alpha$ before critical field long-range effects are pronounced, leading to sharper transitions in $\Pi/N$ and small chains are more effective in these regimes, showing larger values of $\Pi/N$. A higher thermal gradient enhances the sensitivity of $\Pi/N$ to variations in $h_{i}$. For example, in the third figure, the large difference between $\beta_{c}=2.0$ and $\beta_{h}=0.6$ causes a significant shift in $\Pi/N$ post-transition.faor a small thermal gap (e.g., $\beta_{c}=1.0,\beta_{h}=0.5$ in the second figure) result in a smoother decline of $\Pi/N$, with less abrupt changes across the critical point but with a lower performance than the others.
Before the critical point ($h_{i}<1$), small chains exhibit higher $\Pi/N$, reflecting their ability to efficiently capture long-range correlation in weak field $h_{i}$ , while large chains have a lower $\Pi/N$ but show a better performance than the shorter chain for higher $h_{i}$ special in the vicinity of critical point and after phase transition, indicating the growing influence of long-range correlations after the critical point ($h_{i}>1$), large chains outperform small chains, with a higher $\Pi/N$ due to the dominance of long-range correlations. Small chains show a significant decline in $\Pi/N$, reflecting reduced efficiency in capturing long-range effects. At the critical point ($h_{i}=1$), $\Pi$ peaks for large chains as the system approaches Carnot efficiency, reflecting enhanced work output and system efficiency. For small chains, the peak is less pronounced, indicating their limited capacity to utilize critical phenomena effectively.
Refer to captionRefer to caption
; Refer to captionRefer to caption
Figure 14: The evolution of the maxima as a function of $N$ (the number of particles) for the diamagnetic peak (in red) and the paramagnetic peak (in blue) of $\Pi/N$ and $\Pi_{R}/N$, with a comparison to the critical field (in red).
For the case of short-range correlations, we observe an underperformance after the phase transition, but with a stronger dependence on system size, as longer chains prove more efficient both before and after the critical point. It is also noted that for large $\alpha$ values, short-range correlations dominate, and the system’s efficiency is enhanced by the thermal gap, which promotes short thermal correlations and better protects them from thermal fluctuations.
@@
The behavior of $\Pi/N$ as a function of $h_{i}$ demonstrates the intricate relationship between work, system efficiency, and particle number $N$. Large chains tend to dominate in regimes where long-range correlations are significant, especially beyond the critical point, whereas small chains are more effective in the short-range, pre-transition regime. The parameters $\alpha$, $\beta_{h}$, and $\beta_{c}$ play crucial roles in shaping this behavior, with their values determining the system’s sensitivity to field variations and the overall efficiency relative to Carnot. Understanding these dynamics is essential for optimizing quantum heat engines across different operating regimes.

### v2

@@ III Quantum heat engine and Refrigerator
Figure 10: Work per spin $W/N$ and efficiency $\eta$ as functions of the initial transverse field $h_{i}$ for different system size. For quenches performed in the paramagnetic phase, i.e., the critical field is located at $h_{i}=1$ for $\alpha=0.55$, and approximately at $h_{i}\approx 0.75$ for $\alpha=1.2$. for $h_{i}>h_{\text{crit}}$, a distinct structure becomes apparent. In particular, the first peak, which becomes more pronounced as $T_{c}$ increases, exhibits a superextensive scaling of the work with the system size $N$, suggesting a potential influence of quantum critical behavior, same analysis as in the top panels, now applied to the efficiency $\eta$ of the heat engine. The qualitative features remain consistent, including the dependence on NN near the critical peak.
@@ III.1 Quantum Refrigerator
A thermodynamic advantage is expected in the vicinity of the phase transition as highlighted in [57], in finite-temperature systems near criticality, enhanced heat extraction is anticipated around the critical point. This enhancement arises from the divergence of the specific heat [58], which allows substantial heat exchange even under small temperature gradients. A comparable reasoning applies to our magnetic system: as the quantum critical point is approached, the magnetic susceptibility—defined as the derivative of magnetization with respect to the external field—diverges. Consequently, the magnetization becomes highly responsive to even minor variations in the field. Since the work output due to a magnetic field change is proportional to the magnetization [59, 60], this leads us to expect an increase in work extraction in the vicinity of the quantum critical point , similarly, improved heat extraction is expected in the refrigerator mode. Our findings, as depicted in Fig. 8.a-b, demonstrate that for both values of $\alpha$, an increase in the interaction range enhances the strength of spin correlations and mitigates perturbations induced by neighboring particles, particularly in the ferromagnetic phase ($h_{i}<h_{c}$). This fosters coherent spin alignments, thereby improving efficiency. We observe superior heat extraction per spin, $Q_{c}/N$, for long-range interactions, which facilitate more effective energy transfer among nearest neighbors compared to short-range interactions that propagate more locally, thus minimizing $Q_{c}/N$. Long-range interactions lead to a smoother and more gradual decline in $Q_{c}/N$ beyond the critical transition, especially for longer chains, which exhibit greater stability and resilience. Conversely, short-range interactions result in a sharper drop around the critical point, attributable to the localized nature of energy diffusion compared to the more extended correlations observed for smaller $\alpha$. Drawing on the evidence provided by [14], the authors showed that taxonomy separates three regimes by $\alpha$. At $\alpha=0.2$ (very long range, effectively near all-to-all) the system belongs to the non-additive, strong-long-range sector: without Kac rescaling the interaction energy grows super-extensively; collective, mean-field–like correlations and long-lived quasi-stationary states appear, relaxation is anomalously slow, and pronounced finite-$N$ peaks or apparent saturations dominate observables. By contrast, at $\alpha=1.2$ the model lies in the crossover from weak long-range to short-range behaviour, where energy is extensive, critical exponents and finite-size corrections become largely $\alpha$-independent, and observables approach conventional short-range scaling smoothly.
The analysis of the coefficient of performance (COP) corroborates the previously elucidated behavior in Fig. 9.a-b, where an increase in $\alpha$ leads to an enhanced COP, particularly for short chains, which preferentially favor the refrigeration mode. In contrast, long-range interactions, under the influence of neighboring particles, amplify perturbations and spin disorder, a phenomenon that is accentuated in the paramagnetic phase for longer chains, thereby fueling their performance. The contour plots of $Q_{c}/N$ (Fig. 8.a-b) and COP (Fig. 8.c-d) highlight size-dependent dynamics. As $\alpha$ increases, contour lines become densely packed near the transition for small $N$, particularly for long chains, where the behavior is less smooth and more abrupt due to the nature of interactions, which are not as diffusive as long range interaction.
Figure 11: Contour plot of the work per spin $W/N$ and $\eta$ as a function of $h_{i}$ for different sizes N , for different values of interaction range $\alpha$ .
Overall, the observed deviations for small values of $N$ are indeed expected and can be attributed to well-documented finite-size effects, including discretized energy levels, boundary contributions, and the lack of a thermodynamic limit, cause fluctuations and non-extensive behaviors in quantities like heat and work per spin in small spin chains [61, 62]. Quantum coherence and entanglement further influence the performance of quantum thermal machines at small $N$, leading to non-classical thermodynamic signatures. As $N$ increases, these effects diminish, and observables like work and heat per spin align with the extensivity principle of thermodynamics, consistent with known physics of finite-size systems [63, 64]. The system’s behavior strongly depends on the chain size and the correlation regime, with each configuration offering specific advantages depending on the magnetic field. The performance per spin is better in the ferromagnetic phase, with efficiency varying depending on the type of correlation $\alpha$.
@@ III.2 Quantum Heat Engine
We will see in this section the evolution of the work per spin ($W/N$) and efficiency ($\eta$) as a function of the initial transverse field ($h_{i}$) for different values of $\alpha$, $T_{h}$, and $T_{c}$, analyzing the effects of the phase transition and the system’s behavior for small and large chains ($N$). Each graph in Fig.11 highlights how interactions and thermal conditions influence the system’s performance.
Fig 10.a-b elucidates the behavior of the work per spin, $W/N$, as a function of varying $\alpha$, $T_{h}$, and $T_{c}$, highlighting the influence of chain length $N$, transverse field $h_{i}$, and differing interaction ranges. Increasing $\alpha$ reduces the influence of neighboring particles while enhancing localized correlations, thereby promoting stronger spin alignment in the ferromagnetic phase ($h_{i}<h_{c}$). As illustrated in Fig 10-b, elevating $\alpha$ generally leads to improved work extraction per spin, owing to diminished perturbations from proximate neighbors, resulting in a more linear behavior for large $N$. This contrasts with the case of $\alpha=0.55$, where more pronounced perturbative effects from neighboring particles intensify with increasing $N$, leading to a rapid misalignment of spins and consequently higher work extraction for smaller $N$. In contrast, the paramagnetic phase ($h_{i}>h_{c}$), characterized by disordered spins, leverages these neighboring perturbations as a driving force, significantly enhancing work extraction, particularly for longer chains. Regarding the efficiency, $W/Q_{h}$, Figure 10c-d highlights distinct regimes where, in contrast to the work per spin, an increase in $\alpha$ diminishes the cycle’s efficiency. This is attributed to $Q_{h}$, the heat absorbed from the hot reservoir, which is more effectively catalyzed and absorbed under long-range interactions. Conversely, at $\alpha=1.2$, the more localized diffusion imposed by short-range interactions restricts this process, leading to reduced efficiency.Complementing this analysis, Fig. 11.a-b illustrates the contour lines of $W/N$, revealing dependencies on $\alpha$ and the thermal gap. For small $\alpha$, two peaks emerge for short and long chains in phases exhibiting distinct behaviors within the same interaction range, characterized by concentric circles centered around small $N$ in the ferromagnetic phase and large $N$ in the post-critical phase. For larger $\alpha$, a single peak at $\alpha=1.2$ dominates in the post-critical phase for large chains, where the locality of short-range interactions introduces a linear dependence before the phase transition, followed by a gradual decline toward higher field strengths.
Figure 12: The scaling factor per spin $\Pi/N$ ( $\Pi_{R}/N$) as a function of $h_{i}$, for various system sizes NN and different interaction ranges characterized by $\alpha$.
Regarding efficiency, $W/Q_{h}$, Fig. 11.c-d highlights distinct regimes. For $\alpha=0.55$, efficiency decreases gradually, with a rapid vertical drop for long chains before the phase transition, in contrast to $\alpha=1.2$, which exhibits a more tempered decline and varies horizontally due to the linearity of short-range interactions. Thus, the relationship between efficiency and work per spin reflects the complex interplay of thermal fluctuations and correlations in the system, especially across different interaction ranges.
III.3 Scaling factor of the quantum heat engine and refrigerator
The scaling factor $\Pi$ measures how effectively an engine produces work despite having an efficiency lower than the Carnot limit [65]. Furthermore , an elevated $\Pi$ plays a pivotal role in criticality and phase transitions, where it indicates enhanced system stability against critical fluctuations and the capacity to sustain optimal performance during such transitions. Consequently, a high scaling factor reflects increased efficiency, resilience to perturbations, and optimal utilization of energy resources, even though the absolute efficiency remains below the theoretical maximum.To observe the impact and influence of correlations without considering the size of the chain, we will normalize by dividing by $N$ , defined as :
@@ III.3 Scaling factor of the quantum heat engine and refrigerator
where $W$ is the work done by the system, $\eta_{\text{Carnot}}=1-\frac{T_{c}}{T_{h}}$ is the Carnot efficiency, and $N$ is the number of particles in the system in Fig. 12 . These findings underscore the interplay of chain size, field strength, and interaction range in dictating system performance across phase transitions. Physically, $\Pi/N$ reflects the system’s capacity to harness quantum correlations for work output, modulated by the interplay between $\alpha$ and the thermal gradient. Near criticality in finite-temperature systems, enhanced heat extraction is anticipated due to the divergence of specific heat, allowing significant heat exchange with minimal temperature gradients. Similarly, in this magnetic system, the magnetic susceptibility—diverges near the quantum critical point, making magnetization highly responsive to small field changes. Since work output from a magnetic field variation is proportional to magnetization , this heightened sensitivity boosts work extraction at criticality.An increase in $\alpha$ in Fig. 12-b enhances the strength of spin-spin correlations and reduces perturbations induced by neighboring particles, particularly in the ferromagnetic phase ($h_{i}<h_{c}$), promoting coherent spin alignments and elevating $\Pi/N$ by maximizing work extraction. Conversely, lower $\alpha$ values in Fig. 12-a lead to long-range interactions with faster decay, increasing scattering and broadening the phase transition into a wider critical region . This renders short chains more stable before the phase transition. This diminishes the magnitude of $\Pi/N$ in the ferromagnetic phase due to weaker correlations but mitigates the impact of critical fluctuations, resulting in a smoother decline of $\Pi/N$ across $h_{c}$, particularly in smaller systems less sensitive to long-range effects . Let’s move on to the refrigerator mode, we operate with $\Pi_{R}/N$,
$\Pi_{R}/N=\frac{Q_{c}}{(\eta_{\text{RCarnot}}-\frac{Q_{c}}{W})}/N,$ (23)
where $Q_{c}$ denote the heat exchanged during the thermalisation with the cold bath done by the system, $\eta_{\text{RCarnot}}=\frac{1}{T_{h}/T_{c}-1}$ is the Carnot refrigerator efficiency , which reflects the system’s capacity to efficiently extract heat by leveraging spin correlations, modulated by $\alpha$ and the thermal gradient. Similarly to the engine mode, an increase in $\alpha$ influences the strength of correlations with neighboring spins (Fig. 12-c) .
Figure 13: The evolution of the maxima as a function of NN (the number of particles) for the ferromagnetic peak (in red) and the paramagnetic peak (in blue) of $\Pi/N$ and $\Pi_{R}/N$, with a comparison to the critical field (in red).We will also analyze the peaks in the paramagnetic phase before the phase transition and the ferromagnetic phase after the transition, as well as their proportionality and dependence on N for the quantum heat engine and refrigerator mode by using a regression to calculate the exponent $a$ responsible for the growth rate of the scaling factor per spin, proportional to $N^{a}$.
However, unlike the engine mode, heat extraction from the reservoir is more effective when $\alpha$ is lower. This is explained by the divergence of the specific heat near the phase transition, which diminishes as $\alpha$ increases, resulting in a reduced potential for heat extraction and thus providing greater stability for long-range interactions.
@@
In the presence of short-range correlations, post-transition performance declines, with a notable dependence on system size; longer chains demonstrate greater efficiency across all regimes. Additionally, for large $\alpha$, short-range correlations prevail, bolstered by the thermal gap, which mitigates thermal fluctuations and enhances system efficiency. The behavior of $\Pi/N$ as a function of $h_{i}$ demonstrates the intricate relationship between work, system efficiency, and particle number $N$. Large chains tend to dominate in regimes where long-range correlations are significant, especially beyond the critical point, whereas small chains are more effective in the short-range, pre-transition regime.